OpenStudy (anonymous):

1/10,000 times 614.9 times 10^7

OpenStudy (aum):

\[\large \frac{1}{10,000} \times 614.9 \times 10^7 = 10^{-4} \times 614.9 \times 10^7 = \\ \large 614.9 \times 10^{7-4} = 614.9 \times 10^3 = 614900 ~~\text{or } ~~6.149 \times 10^5 \]

OpenStudy (anonymous):

@aum thanks really appreciate it :)

OpenStudy (aum):

ou are welcome.Y